# REDUCIO! Generating 10241024 Video within 16 Seconds using Extremely Compressed Motion Latents

### v2

@@ Abstract
Commercial video generation models have exhibited realistic, high-fidelity results but are still restricted to limited access. One crucial obstacle for large-scale applications is the expensive training and inference cost. In this paper, we argue that videos contain much more redundant information than images, thus can be encoded by very few motion latents based on a content image. Towards this goal, we design an image-conditioned VAE to encode a video to an extremely compressed motion latent space. This magic Reducio charm enables 64 reduction of latents compared to a common 2D VAE, without sacrificing the quality. Training diffusion models on such a compact representation easily allows for generating 1K resolution videos. We then adopt a two-stage video generation paradigm, which performs text-to-image and text-image-to-video sequentially. Extensive experiments show that our Reducio-DiT achieves strong performance in evaluation, though trained with limited GPU resources. More importantly, our method significantly boost the efficiency of video LDMs both in training and inference. We train Reducio-DiT in around 3.2K training hours in total and generate a 16-frame 10241024 video clip within 15.5 seconds on a single A100 GPU. Code released at https://github.com/microsoft/Reducio-VAE.
[width=autoplay=True,keepaspectratio]8figs/vsample_1/000015
@@
[width=autoplay=True,keepaspectratio]8figs/vsample_2/000015
@@ 1 Introduction
The straightforward but surprisingly dangerous charm, Reducio, cause certain things to shrink.
Miranda Goshawk “Harry Potter” [41]
Recent advances in latent diffusion models (LDMs) [40] for video generation have presented inspiring results [59, 4, 1, 3, 51, 62, 2, 5, 36], showing great potential in various applications. Commercial models like Sora [31], Runway Gen-3 [42], Movie Gen [35], and Kling [21] can already generate photorealistic and high-resolution video clips. However, training and deploying such models are computationally demanding—thousands of GPUs and millions of GPU hours are required for training, and the inference for a one-second clip costs several minutes. Such a high cost is becoming a significant obstacle for research and large-scale real-world applications.
@@
More importantly, as the image prior delivers enriched spatial information of video contents, factorized video generation offers the potential for further compression. We argue that video can be encoded into very few latents representing motion variables, i.e., motion latents, plus a content image. Hereby, we introduce the magical Reducio charm by building a VAE comprised of a 3D encoder that aggressively compresses input videos into a down-sampled compact space, and a 3D decoder fused with feature pyramids of the middle frame as the content condition. Notably, the resulting Reducio-VAE surpasses the common 2D VAE by 5db in PSNR while using a smaller latent space.
@@
In summary, our contributions are as follows: (1) We encode video into very few motion latents plus a content image, where the designed Reducio-VAE can compress video into a smaller latent space compared to common 2D VAE. (2) Based on Reducio-VAE, a diffusion model called Reducio-DiT is devised by incorporating additional image conditions. (3) Experiments show that our Reducio-DiT can significantly speed up the generation process while producing high-quality videos.
@@ 2 Related Work
Video diffusion models have become the latest craze in the field of text-to-video generation. Due to the huge success of text-to-image LDMs [14, 28, 44, 27, 40], attempts [4, 3, 51] have been made to leverage the powerful UNet-based stable diffusion models pre-trained in 2D space by inserting temporal convolution and attention layers. Recently, the strong performance of transformers has continued in image generation domain [8, 6, 22] by diffusion transformers (DiT) [32]. Consequently, recent works have focused on extending DiT-based diffusion models from 2D spatial attention to spatiotemporal attention.
While the mainstream exploration [35] follows the paradigm of image diffusion models by using text as the only condition, an emerging stream of works [10, 52, 65] decompose text-to-video diffusion into two steps, i.e., text-to-image and text-image-to-video generation. This factorizing encourages models to concentrate on motion modeling instead of challenging spatiotemporal joint modeling. In this work, we take advantage of factorized text-to-video generation and build our framework upon DiT. Thanks to the delicate structure of Reducio-VAE, our model manages to preserve the fine details of image priors in generated videos, and hence can be used for more high-demanding real-world applications, e.g., advertisement generation.
Efficient diffusion models have gained increasing attention due to the tremendous computational resources and time required for training and inferring samples, especially for high-resolution images and videos. On one hand, researchers have made attempts to build LDMs with more efficient architectures [9, 48, 56, 69]. For example, DIM [48] uses Mamba as its backbone to ease the computational burden compared to vanilla transformers. Another line of work leverages efficient diffusion techniques to perform fewer sampling steps during inference [25, 45, 61] or to accelerate convergence during training [11, 55]. For instance, Stable Diffusion 3 [8] and Flux [22] employ rectified flow [23] for improved training and inference efficiency.
@@ 3.1 Reducio Video Autoencoder
Reducio charm on videos. As adjacent video frames share large similarities in pixel space, video signals are inherently more redundant than images. Based on this idea, we adopt a more aggressive down-sampling factor for video autoencoder. Specifically, we inflate the 2D convolutions into 3D convolutions and increase the number of autoencoder blocks, increasing and to 4 and 16, as illustrated in (b) of Fig. 2. While the overall down-sampling factor increases by , the compressed latent space suffices to achieve comparable pixel-space reconstruction performance, making a strong case for the potential of extremely shrinking video latent dimension.
As for image-to-video generation, we argue that video latent can be further compressed as the content image contains rich visual details and substantial overlapping information. In such a scenario, video latents can be further simplified to represent the motion information. We can find similar spirits in video codec [39], where carefully designed modules use reference frames as inputs to perform high-fidelity reconstruction of a given compressed code. Specifically, we propose to cast the Reducio charm on video latent space, as shown in (c) of Fig. 2. Generally, we choose the middle frame as the content guidance, compress input videos by an extreme down-sampling factor with a 3D encoder, and reconstruct from the latent space with the aid of content frame features pyramids. Specifically, we infuse content features of sizes varying in and via cross-attention within the decoder, adopt an extreme compression ratio, i.e., and . While Reducio-VAE projects videos into more compressed latent space than one of SDXL-VAE [34], it significantly outperforms the latter in reconstruction performance.
@@ 3.2 Reducio Diffusion Transformer
Thanks to the magical compression capability of Reducio-VAE, we project input videos into compressed latent space with a down-sampling factor of 4096, and hence significantly speed up the training and inference of diffusion models. In specific, we adopt the DiT-XL [32] model and follow most of the design choices adopted by PixArt- [6], i.e., AdaLN-single modules, cross attention layers with Flan-T5-XXL [37] text conditions. The training target follows the reverse diffusion process [14], which can be depicted as
where is latent, is time step, and indicate the mean and variance of the sample at the current time step. By re-parameterizing as a noise prediction network , the model can be trained using simple mean-squared error between the predicted noise and the ground truth sampled Gaussian noise formulated as
To adapt the image diffusion model to video, we consider two options: (1) Directly transforming the 2D attention to full 3D attention, without adding additional parameters. (2) Adding temporal layers and transforming the model to perform 2D spatial attention plus 1D temporal attention. We adopt option (1) by default. Studies on the evaluation of these two options are shown in Sec. 4.3.2.
In addition, to tailor for image-conditioned video generation, we introduce Reducio-DiT (illustrated in Fig. 3) with additional image-condition modules, as described below.
Content frame modules consist of a semantic encoder built upon pretrained OpenCLIP ViT-H [7] and a content encoder initialized with the SD2.1-VAE [40]. The former encoder projects content frames into a high-level semantic space while the latter one mainly focuses on extracting the spatial information. We concatenate the obtained features with text tokens output by T5 to form the image-text-joint condition, and then cross-attend them with the noisy video latent. Concretely, the detailed operations can be described as:
where is the condition image, refers to element-wise addition, refers to concatenation, and refers to cross-attention.
@@
To better accommodate the model to videos with varying aspect ratios, we inject the aspect ratio and size of input videos as embeddings into the DiT model. Then we add the dynamic size embedding with timestep embedding before being fed into the MLP. More importantly, for models in the first and the second stages, the content encoder cross-attends tokens with the noisy latent. As computational costs grow rapidly when resolution increases, we introduce a strategy to ensure the efficiency of Reducio-DiT on high-resolution videos. Specifically, we aims to shrink the content embedding by in the spatial dimension in each cross-attention layer. Inspired by Deepstack [26], we divide input tokens into 4 groups of grids and iteratively concatenate each group of tokens with the other condition tokens.
@@ 4.1 Training and Evaluation Details
Our Reducio-VAE is trained on a 400K video dataset collected from Pexels***https://www.pexels.com/. Pexels contains a large number of free and high-quality stock videos, where each video is accompanied by a short text description. We train Reducio-VAE from scratch on videos with 16 FPS. Note that we split videos with higher resolutions into tiles in the spatial dimension. To perform the VAE comparison, PSNR, SSIM [54], LPIPS [67] and reconstruction FVD (rFVD) [49] are employed as the evaluation metrics.
@@
A child excitedly swings on a rusty swing set, laughter filling the air.
We adopt DPM-Solver++ [24] as the sampling algorithm for efficient inference and set the sampling step to 20. For evaluation in Tab. 2, we calculate the speed of generating 16-frame video clips on a single A100 80G GPU and denote out-of-memory as OOM. For Reducio-DiT, we use PixArt--256 and PixArt--1024 to generate the content frame for and videos, respectively. We report the FVD and IS scores of Reducio-DiT-512 under the zero-shot setting on UCF-101 [46] and MSR-VTT [60]. Since our method is a two-stage video generation approach, which requires a condition image prior, we follow MicroCinema [52] to use SDXL [34] to generate the condition image for all the evaluations, including the FVD on UCF-101 and MSR-VTT. Additionally, we evaluate Reducio-DiT-512 on the recent video generation benchmark VBench [17].
@@ 4.2 Main Results
We observe Reducio-VAE has worse rFVD than others on UCF-101. By careful inspection, we find the selected prior content image is blurring in many cases due to the low visual quality of UCF-101 and hence result in a lower reconstruction quality. In contrast, Reducio-VAE places first on rFVD on pexels, which consists of videos with much improved visual quality and much higher resolution.
Reducio-DiT achieves a promising balance between efficiency and performance. As we harness an extremely compressed latent space, the training cost of our diffusion models is similar to the one of the text-to-image diffusion model. With a training cost of merely 3.2K A100 hours, our model achieves FVD scores of 318.50 on UCF-101 and 291.0 on MSR-VTT and the VBench score of 81.39, beating a range of previous stat-of-the-art video LDMs, as recorded in Tab. 2. Fig. 5 displays the comparison between Reducio-DiT and the other two stat-of-the-art image-to-video models. While DynamicCrafter and SVD-XT both fail to generate stable frames consistent with the given content frame, Reducio-DiT yields high visual quality with facial details in the content frame uncorrupted.
Moreover, Reducio-DiT significantly boosts the throughput of video LDMs. The compressed latent space facilitates Reducio-DiT to output videos of 16 frames in almost 1 second, which realize a speed up compared with Lavie. Moreover, 16-frame video generation costs Reducio-DiT only for 15.4 seconds on average, leading to a striking speedup of over Lavie. We argue the core philosophy of Reducio lies in the extremely compressed latent space, which acts as the key role in accelerating both the training and inference speed of video LDMs.
@@ 4.3.1 Design of Reducio-VAE
Spatial redundancy is the key for compression. As Tab. 3 displays, when adopting an identical overall down-sampling factor , Reducio-VAE achieves worse performance with larger compression factor in the temporal dimension. The result suggests that when a reference frame is provided, videos contain heavier redundancy in the spatial dimension than in the temporal dimension. Similarly, when is kept unchanged, increasing from 2 to 4 leads to a drop in PSNR by 0.91 db. To balance the trade-off between more compact latent space and better reconstruction performance, we set and in default.
@@ 4.3.2 Design of Reducio-DiT
Using joint spatiotemporal attention (3D attention) outweighs using factorized spatial and temporal attention (i.e., 2D + 1D attention) in generation quality. Interestingly, we observe that factorized attention leads to a faster convergence of training loss. However, with the same training step, as shown in Tab. 7, factorized attention lags behind its counterpart with joint 3D attention for 45 in FVD. We suppose the possible reason is that 2D + 1D scheme demands adding additional temporal layers and performs factorized self-attention on a small set of tokens e ach, making it hard to model smooth open-set motion with the light computation. In contrast, 3D attention directly exploits the original parameters and collaborates all spatiotemporal tokens.

### v3

# REDUCIO! Generating 1K Video within 16 Seconds using Extremely Compressed Motion Latents

@@ Abstract
Commercial video generation models have exhibited realistic, high-fidelity results but are still restricted to limited access. One crucial obstacle for large-scale applications is the expensive training and inference cost. In this paper, we argue that videos contain significantly more redundant information than images, allowing them to be encoded with very few motion latents. Towards this goal, we design an image-conditioned VAE that projects videos into extremely compressed latent space and decode them based on content images. This magic Reducio charm enables 64 reduction of latents compared to a common 2D VAE, without sacrificing the quality. Building upon Reducio-VAE, we can train diffusion models for high-resolution video generation efficiently. Specifically, we adopt a two-stage generation paradigm, first generating a condition image via text-to-image generation, followed by text-image-to-video generation with the proposed Reducio-DiT. Extensive experiments show that our model achieves strong performance in evaluation. More importantly, our method significantly boosts the training and inference efficiency of video LDMs. Reducio-DiT is trained in just 3.2K A100 GPU hours in total and can generate a 16-frame 10241024 video clip within 15.5 seconds on a single A100 GPU. Code is available at https://github.com/microsoft/Reducio-VAE.
@@ 1 Introduction
Recent advances in latent diffusion models (LDMs) [45] for video generation have presented inspiring results [71, 4, 1, 3, 61, 74, 2, 5, 41, 67, 56, 53, 55, 54], showing great potential in various applications. Commercial models like Sora [36], Runway Gen-3 [46], Movie Gen [40], and Kling [24] can already generate photorealistic and high-resolution video clips. However, training and deploying such models are computationally demanding—thousands of GPUs and millions of GPU hours are required for training, and the inference for a one-second clip costs several minutes. Such a high cost is becoming a significant obstacle for research and large-scale real-world applications.
@@
More importantly, as the image prior delivers enriched spatial information of video contents, factorized video generation offers the potential for further compression. We argue that video can be encoded into very few latents that represent motion variables, i.e., motion latents, plus a content image. Hereby, we introduce the magical Reducio charm by building a VAE comprised of a 3D encoder that aggressively compresses input videos into a down-sampled compact space, and a 3D decoder fused with feature pyramids of the middle frame as the content condition. Notably, the resulting Reducio-VAE surpasses the common 2D VAE by 5db in PSNR while using a smaller latent space.
@@
In summary, our contributions are as follows: (1) We encode video into very few motion latents plus a content image, where the designed Reducio-VAE can compress video into a smaller latent space compared to common 2D VAE. (2) Based on Reducio-VAE, a diffusion model called Reducio-DiT is devised by incorporating multimodal conditions to perform text-image to video generation. (3) Experiments show that our Reducio-DiT can significantly speed up the generation process while producing high-quality videos.
@@ 2 Related Work
Video diffusion models have become the latest craze in the field of text-to-video generation. Due to the huge success of text-to-image LDMs [17, 34, 48, 33, 45], attempts have been made to leverage the powerful UNet-based stable diffusion models pre-trained in 2D space by inserting temporal convolution and attention layers [12, 70, 4, 3, 61]. Recently, the strong performance of transformers has continued in image generation domain [60, 31, 10, 6, 25]. Consequently, recent works have focused on extending 2D diffusion transformer (DiT) [37] to spatiotemporal domain.
While the mainstream exploration [40, 85, 26] follows the paradigm of image diffusion models by using text as the only condition, an emerging stream of works [13, 63, 78] decompose text-to-video diffusion into two steps, i.e., text-to-image and text-image-to-video generation. The factorized framework encourages models to focus on motion modeling instead of challenging spatiotemporal joint modeling. In this work, we build an image-to-video model upon DiT. Thanks to the delicate design of Reducio-VAE, our diffusion model preserves fine details of image priors in generated videos and hence can be used for more high-demanding real-world applications, e.g., advertisement generation.
Efficient diffusion models have gained increasing attention due to the tremendous computational resources and time required for training and inferring samples, especially for high-resolution images and videos. On one hand, researchers have made attempts to build LDMs with more efficient architectures [11, 52, 68, 84]. For example, DIM [52] uses Mamba as its backbone to ease the computational burden compared to vanilla transformers. Another line of work leverages efficient diffusion techniques to perform fewer sampling steps during inference [28, 30, 49, 73, 80] or to accelerate convergence during training [14, 66]. For instance, CausVid [75] applies distribution matching distillation between a causal student video generation model and a bidirectional teacher model.
@@ 3.1 Reducio Video Autoencoder
Reducio charm on videos. As adjacent frames share large similarities in pixel space, video signals inherently contain more redundancy than images. Therefore, we apply a more aggressive down-sampling strategy for video autoencoder. Specifically, we employ 3D convolutions and increase the number of autoencoder blocks, increasing and to 4 and 16, as shown in (b) of Fig. 2, respectively. Though the overall down-sampling factor increases by , the compressed latent space suffices to achieve comparable pixel-space reconstruction quality, highlighting the potential for extremely shrinking video latent dimensions.
As for image-to-video generation, we argue that video latent can be further compressed as the content image contains rich visual details and substantial overlapping information. In such a scenario, video latents can be further compressed to represent the motion information. We can find similar spirits in the video codec [44], where carefully designed modules use reference frames as inputs to perform high-fidelity reconstruction of a given compressed code. Specifically, we propose to cast the Reducio charm on video latent space, as shown in (c) of Fig. 2. Generally, we choose the middle frame as the content guidance, compress input videos by an extreme down-sampling factor with a 3D encoder, and reconstruct from the latent space with the aid of content frame features pyramids. Specifically, we infuse content features of sizes varying in and via cross-attention within the decoder, adopt an extreme compression ratio, i.e., and . While Reducio-VAE projects videos into more compressed latent space than one of SDXL-VAE [39], it significantly outperforms the latter in reconstruction performance.
Note that we train Reducio-VAE on videos of resolution. To encode and decode high-resolution videos without exceeding the memory limit, we split the input video into overlapping tiles along the spatial dimension, following the approach used in Movie Gen [40]. In practice, we use a tile size of 256 (8 in latent space) with an overlap of 64 pixels (2 latents). Overlapping latents and decoded video regions are blended using linear combinations.
@@ 3.2 Reducio Diffusion Transformer
Thanks to the magical compression capability of Reducio-VAE, we project input videos into compressed latent space with a down-sampling factor of 4096, hence significantly speeding up the training and inference of diffusion models. Specifically, we adopt the DiT-XL [37] model and follow most of the design choices adopted by PixArt- [6], i.e., AdaLN-single modules, cross attention layers with Flan-T5-XXL [42] text conditions.
To adapt the image diffusion model to video, we consider two options: (1) Directly transforming the 2D attention to full 3D attention without adding additional parameters [74, 8]. (2) Adding temporal layers and transforming the model to perform 2D spatial attention plus 1D temporal attention [85, 78, 26]. We adopt option (1) by default. Studies on the evaluation of these two options are shown in Sec. 4.3.2. In addition, we introduce Reducio-DiT (illustrated in Fig. 3) with additional image-condition modules, as described below.
Content frame modules consist of a semantic encoder built upon pretrained OpenCLIP ViT-H [9] and a content encoder initialized with the SD2.1-VAE [45]. The former encoder projects content frames into a high-level semantic space while the latter one mainly focuses on extracting the spatial information. We concatenate the obtained features with text tokens output by T5 to form the image-text-joint condition, and then cross-attend them with the noisy video latent.
@@
To better accommodate the model to videos with varying aspect ratios, we inject the aspect ratio and size of input videos as embeddings [6] into the DiT model. We then add the dynamic size embedding with timestep embedding before being fed into the MLP. More importantly, for models in the first and the second stages, the content encoder cross-attends tokens with the noisy latent. As computational costs grow rapidly when resolution increases, we introduce a strategy to ensure the efficiency of Reducio-DiT on high-resolution videos. Specifically, we aim to shrink the content embedding by in the spatial dimension in each cross-attention layer. Inspired by Deepstack [32], we divide input tokens into 4 groups of grids and iteratively concatenate each group of tokens with the other condition tokens.
@@ 4.1 Training and Evaluation Details
Our Reducio-VAE is trained on a 400K video dataset collected from Pexels***https://www.pexels.com/. Pexels contains a large number of free and high-quality stock videos, where each video is accompanied by a short text description. We train Reducio-VAE from scratch on videos with 16 FPS. To perform the VAE comparison, PSNR, SSIM [65], LPIPS [81] and reconstruction FVD (rFVD) [57] are employed as the evaluation metrics.
@@
| Model | FVD | |
| UCF-101 | MSR-VTT | |
| VideoComposer [62] | 576.8 | 377.3 |
| I2VGen-XL [83] | 571.1 | - |
| DynamiCrafter [69] | 429.2 | 234.7 |
| Reducio-DiT | 318.5 | 291.9 |
We adopt DPM-Solver++ [29] as the sampling algorithm for efficient inference and set the sampling step to 20. For evaluation in Tab. 3, we use PixArt--256 and PixArt--1024 to generate the content frame for and videos, respectively. We denote out-of-memory as OOM and calculate the speed of generating 16-frame video clips on a single A100 80G GPU, including both text-to-image and text-image to video cost. For Reducio-DiT, We report the FVD and IS scores of Reducio-DiT-512 under the zero-shot setting on UCF-101 [50] and MSR-VTT [72]. Since our model performs image-to-video generation and requires a condition image prior for the text-to-video pipeline, we follow MicroCinema [63] to use SDXL [39] to generate the condition image and LLaVA-1.5 [27] to generate captions for FVD evaluations on UCF-101 and MSR-VTT. Additionally, we evaluate Reducio-DiT-512 on the recent video generation benchmark VBench [20].
@@ 4.2 Main Results
We observe that Reducio-VAE has worse rFVD than others on UCF-101. By careful inspection, we find the selected prior content image is blurring in many cases due to the low visual quality of UCF-101 and hence results in a lower reconstruction quality. In contrast, Reducio-VAE places first on rFVD on pexels, which consists of videos with much improved visual quality and much higher resolution.
Reducio-DiT is a powerful image-to-video model. As shown in Tab. 2, Reducio-DiT achieves a 90-point lower FVD in the zero-shot setting on UCF-101 compared to the state-of-the-art image-to-video model, DynamiCrafter. Additionally, Reducio-DiT surpasses VideoComposer in FVD metrics on MSR-VTT but falls short of DynamiCrafter in this aspect. For qualitative evaluation, we conduct a user study using identical images generated by PixArt--1024 as input. Participants evaluate 15 video groups based on three criteria: (1) visual quality, (2) motion quality, and (3) alignment of the condition image. Responses from 20 participants are presented in Fig. 6.
Notably, Reducio-DiT significantly outperforms in visual quality (64%) and content alignment (57%) by overcoming temporal flickering and preserving fine details from the condition image. Moreover, thanks to its strong temporal consistency, Reducio-DiT achieves a preference rate comparable to SVD-XT in motion quality, yet needs much fewer training hours and data. Fig. 6 displays the comparison between Reducio-DiT and the other two stat-of-the-art image-to-video models. While DynamicCrafter and SVD-XT fail to generate stable frames consistent with the given content frame, Reducio-DiT yields high visual quality with maintained facial details. Overall, Reducio-DiT demonstrates strong performance in image-to-video generation.
top: pouring milk into a cup of coffee. bottom: A storm trooper vacuuming the beach.
Reducio-DiT effectively balances efficiency and performance. By integrating Reducio-DiT with a text-to-image LDM for condition image generation, the factorized text-to-video task can be performed efficiently. Consequently, our model achieves FVD scores of 318.5 on UCF-101 and 291.0 on MSR-VTT and the VBench score of 81.39, beating a range of state-of-the-art video LDMs, as recorded in Tab. 3. More importantly, as most mainstream text-to-video models are built upon pre-trained text-to-image LDMs, we compare Reducio-DiT with baselines on training costs for video samples. Thanks to the extremely compressed latent space, we achieve competitive performance with a training cost of merely 3.2K A100 hours.
Moreover, Reducio-DiT significantly boosts the throughput of video LDMs. The compressed latent space facilitates Reducio-DiT to output videos of 16 frames in almost 1 second, which realizes a speed-up compared with Lavie. Moreover, 16-frame video generation costs Reducio-DiT only for 15.4 seconds on average, leading to a striking speed-up of over Lavie. We argue that the core philosophy of Reducio lies in the extremely compressed latent space, which accelerates both the training and inference speed of text-to-video generation.
Although Reducio-DiT is trained exclusively on 16-frame videos sampled at 16 fps, we present that it can be adapted for longer video generation in a zero-shot manner. This is achieved by recursively using the last frame of a generated clip as the input condition for the subsequent clip. Fig. 7 presents the visualization of frames sampled uniformly from 32-frame and 48-frame videos, demonstrating that Reducio-DiT maintains consistent motion while preserving high visual quality.
@@ 4.3.1 Design of Reducio-VAE
Spatial redundancy is the key for compression. As Tab. 4 displays, when adopting an identical overall down-sampling factor , Reducio-VAE achieves worse performance with a larger compression factor in the temporal dimension. The result suggests that when a reference frame is provided, videos contain heavier redundancy in the spatial dimension than in the temporal dimension. Similarly, when is kept unchanged, increasing from 2 to 4 leads to a drop in PSNR by 0.91 db. To balance the trade-off between more compact latent space and better reconstruction performance, we set and in default.